Solve algebraically using only one variable: Find three consecutive odd integers such that three
times the sum of the second two is three less than nine times the first one.

Answer by Boreal(15235)

x, x+2, x+4 are the integers
3(x+2+x+4)+3=9x
3(2x+6)+39x
6x+21=9x
3x=21
x=7
7,9,11
3 times (9+11)+3=9*7
63=63